# Towards Fair RAG: On the Impact of Fair Ranking in Retrieval-Augmented Generation

### v3

@@ Abstract.
Modern language models frequently include retrieval components to improve their outputs, giving rise to a growing number of retrieval-augmented generation (RAG) systems. Yet, most existing work in RAG has underemphasized fair ranking techniques and neglected the diverse interests of all stakeholders. In this paper, we present the first comprehensive study of RAG systems that incorporate fairness-aware rankings, focusing on both ranking fairness and attribution fairness—ensuring equitable exposure of sources cited in the final text. We specifically examine item-side fairness, i.e., whether retrieved documents receive balanced exposure, and assess how this affects both the system’s overall performance and the eventual distribution of cited sources. Across twelve RAG models and seven tasks, we find that fairness-aware retrieval frequently retains or even improves ranking effectiveness and generation quality, countering the widespread belief that fairness compromises system performance. Moreover, we show that fair retrieval leads to more balanced attribution in the final responses, ensuring that the cited sources are credited more equitably. Our results underscore the importance of item-side fairness throughout both retrieval and generation phases, offering key insights for building more responsible and equitable RAG systems and illustrating promising avenues for future exploration in fair ranking and source attribution.  [^1] We publicly release our codebase and dataset at https://github.com/kimdanny/Fair-RAG.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2025/06ccs: Computing methodologies Natural language generationccs: Information systems Evaluation of retrieval results
@@ 3.1. Construction of a Test Collection with Utility Labels
We define an item’s worthiness by the marginal gain in utility (utility-gain) it provides to a language model (specifically, the generator in a RAG system) when used to solve a specific task as part of the augmentation process. To assess this utility-gain, each item in the corpus is individually supplied to the generator along with an input question. The utility-gain is then calculated as the difference between the utility of the augmented generator and that of a baseline language model without any information about the item. Formally, let $u_{i}$ denote the baseline string utility score from the vanilla language model prompted only with the input question, and let $u_{j}$ represent the utility score from the language model with a prompt augmented by the $j$’th item $d_{j}$ in the corpus. The item $d_{j}$ is considered useful if the utility-gain $\delta_{j}=u_{j}-u_{i}$ is positive, and not useful otherwise (see Appendix B).
@@ 3.3.1. Expected Exposure in the Context of Machine Users
Expected Exposure (EE) (Diaz et al., 2020) works by estimating the exposure of items across rankings (e.g., $\sigma$) created by a subject model, and comparing them with an optimal set of rankings that always satisfy the item-fairness. To represent the attention over $n$ items given by the consumer (generator in RAG), an $n\times 1$ system exposure vector $\mathbf{\epsilon}$ is created. This is then compared with an $n\times 1$ target exposure vector $\mathbf{\epsilon}^{*}$, where it represents the exposure of items allocated by an oracle retriever that always rank useful items above non-useful ones (Diaz et al., 2020).
@@ 3.3.3. Expected Attribution Rate
where $\text{NLI}(d,\hat{y}_{\pi})\!=\!1$ if item $d$ entails the output $\hat{y}_{\pi}$, and 00 otherwise, by a natural language inference model. 444 This metric is similar to the context utilization metric from Ru et al. (2024), which also measures how effectively the generator leverages retrieved context. However, Ru et al. (2024) decompose the generated response into multiple claims and focus only on the proportion of relevant items, whereas our approach applies a single NLI check for each retrieved item.
@@ 3.3.5. Normalization of Metrics
Proofs and details on how each metric is normalized by its lower and upper bounds can be found in the Appendix C.
@@ 4. Experiments
We choose the LaMP benchmark (Salemi et al., 2024b) for our dataset. It assesses the personalization capability of language models through retrieval-augmentation of users’ interaction history in a platform. LaMP includes various prediction tasks, such as classification, regression, and generation, and is well-suited for tasks where multiple items can be relevant/useful, unlike QA tasks with typically one or two provenance items. The retrieval items in LaMP have clear providers and consumers, aligning with our goal to ensure fairness for individual item providers. For example, in LaMP-1, retrieval items are academic papers, where exposure can increase citation counts for authors. In LaMP-4, retrieval items are news articles, where exposure can lead to monetary compensation for journalists.  [^5] Throughout this paper, we use LaMP-1 to represent classification and LaMP-4 to represent text generation in visualizations. All LaMP tasks were used in our experiments, and similar trends were observed. Due to the absence of a test set, we constructed a test collection as described in §3.1, using the first thousand entries of a user-based development set. Then, we discarded entries that have only one useful item in the corpus, as it is unnecessary to concern item-fairness in that case. We release the test collection along with dataset statistics. Dataset statistics can be found in the Appendix E.
@@
Figure 3. Effect of a fairness control parameter ($\alpha$) on the disparity of rankings ($\overline{\text{EE-D}}$) in LaMP task 1 and 4. The RAG model is configured with the Contriever and Flan-UL2. Each data point represents the normalized EE-D of each run of the experiment (i.e., one query –¿$N$ sampled rankings –¿ $\overline{\text{EE-D}}$ of the $N$ rankings).
@@ 5. Results
Table 2. Each value in the table is the difference between the utility of a baseline (deterministic) RAG model and the average utility of a fairer RAG model at a specific retrieval disparity interval. Nonnegative differences are highlighted. All LaMP tasks shows similar trend. A superscript * denotes a statistically significant difference between the treatment performance and the baseline performance based on an unpaired Student’s t-test ($p<0.05$). Full results are listed in Appendix D.
@@
Figure 5. Major relationships between retrieval fairness (RetDisparity), ranking quality, expected attribution rate, and consumption fairness (AttDisparity). We empirically show that there is a tradeoff between retrieval fairness and expected attribution rate as there is a positive relationship between ranking quality and attribution rate. Most importantly, we show that retrieval fairness does not necessarily directly propagates to the consumption fairness due to varying attribution rate of a generator. The same interpolation methods are used as Figure 4 and runs were averaged across generators. RetDisparity refers to the disparity among retrieved items, while AttDisparity refers to the disparity among attributed sources.
@@ 7. Conclusion
In future work, we hope to extend this framework to consider graded or missing judgments and exploring the different notions of fairness in RAG systems, ultimately advancing the field of trustworthy RAG systems research.
@@ Appendix A Notation
Appendix B Labeling Procedure
@@ C.1. Normalization of EE-D
For upper bound, recall that the $\mathbf{\epsilon}$ is computed based on samples of rankings from a stochastic policy $\pi$. For relevance-independent browsing models (e.g., MU), all rankings $\pi\in S_{n}$ have identical exposure norms ${\left\lVert\mathbf{\epsilon}^{\pi}\right\rVert}_{2}^{2}$, where $\mathbf{\epsilon}^{\pi}$ is the exposure of items for ranking sampled from $\pi$. Then,
@@
With MU, the upper bound of EE-D, ${\left\lVert\bar{\mathbf{\epsilon}}\right\rVert}_{2}^{2}$ becomes $\sum_{i=1}^{n}{\text{MU}(i)}^{2}=k$. Therefore, per query $q=\phi_{q}(x)$, we calculate a normalized EE-D
(16) $\overline{\text{EE-D}}={\left\lVert\mathbf{\epsilon}\right\rVert}_{2}^{2}/k\quad\in[0,1]$
The same normalization process can be used for EAE-D, but with attributed items.
@@ C.2. Normalization of EE-R
With MU, the upper bound of EE-R can be calculated by equation 6.
@@
(28) $\overline{\text{EE-R}}=\begin{cases}\langle\mathbf{\epsilon},\mathbf{\epsilon}^{*}\rangle/(m+\frac{(k-m)^{2}}{n-m})&\text{($m\leq k$)}\\ \langle\mathbf{\epsilon},\mathbf{\epsilon}^{*}\rangle/(\frac{k^{2}}{m})&\text{($m>k$)}\end{cases}\in[0,1]$
@@ C.3. Normalization of EU
Recall that the string utility $u_{\pi}=\mu_{u}(y,\mathcal{G}(\phi_{p}(x,\pi)))$ is a utility of a RAG model with a sampled ranking $\pi\in\sigma$. Let $\sigma_{\alpha}$ denote a sample of rankings with the fairness control parameter set to $\alpha$. Also, let $\sigma^{*}$ denote the set of sampled permutations (rankings) from the oracle stochastic retriever, a policy that always places relevant items above non-relevant items; thus the oracle generates $m!(n-m)!$ number of unique optimal permutations.
To obtain an approximated upper bound of the utility $u_{max}$, when the runs of the experiments were run with $\alpha=(1,2,4,8)$, we take the maximum over all samples,
(29) $u_{max}=max\left(\{u_{\pi}\}_{\pi\in\sigma_{1}}\cup\{u_{\pi}\}_{\pi\in\sigma_{2}}\cup\{u_{\pi}\}_{\pi\in\sigma_{4}}\cup\{u_{\pi}\}_{\pi\in\sigma_{8}}\cup\{u_{\pi}\}_{\pi\in\sigma^{*}}\right)$
With $u_{max}$, we max-normalize the EU. Since $\frac{1}{N}\sum_{\pi\in\sigma}\frac{u_{\pi}}{u_{max}}$ is the same as $(\frac{1}{N}\sum_{\pi\in\sigma}u_{\pi})/u_{max}$, per query $q$, we get a normalized EU
(30) $\overline{\text{EU}}=\frac{\text{EU}}{u_{max}}\quad\in[0,1]$
@@
Appendix D Fairer RAG Models Vs. Deterministic RAG Models
Disparity Intervals
BM25+FlanT5Small (0.308) -0.12 -0.13 -0.18 -0.02 -0.15
BM25+FlanT5Base (0.670) -0.20 -0.04 -0.08 -0.05 -0.02
BM25+FlanT5XXL (0.531) -0.07 +0.03 +0.02 +0.06 +0.11
BM25+FlanUL2 (0.557) -0.1 -0.02 -0.02 +0.01 +0.03
SPLADE+FlanT5Small (0.241) -0.03 -0.22 +0.19 -0.04 +0.14
SPLADE+FlanT5Base (0.646) -0.15 +0.06 +0.08 0.00 +0.03
SPLADE+FlanT5XXL (0.671) -0.18 -0.16 +0.05 +0.02 +0.01
SPLADE+FlanUL2 (0.632) -0.15 -0.08 +0.02 +0.02 +0.01
Contriever+FlanT5Small (0.286) -0.08 -0.29 -0.06 +0.03 -0.14
Contriever+FlanT5Base (0.637) -0.16 +0.05 -0.06 +0.03 0.00
Contriever+FlanT5XXL (0.651) -0.19 -0.04 -0.11 +0.03 0.00
Contriever+FlanUL2 (0.620) -0.16 -0.09 -0.05 +0.02 -0.03
Table 4. LaMP-1
Disparity Intervals
BM25+FlanT5Small (0.274) -0.13 +0.02 -0.04 -0.04 +0.03
BM25+FlanT5Base (0.223) -0.01 +0.05 +0.04 +0.03 +0.13
BM25+FlanT5XXL (0.310) +0.05 +0.02 +0.24 +0.13 +0.18
BM25+FlanUL2 (0.433) -0.03 +0.07 +0.11 -0.02 +0.09
SPLADE+FlanT5Small (0.209) -0.06 +0.10 +0.06 +0.05 +0.14
SPLADE+FlanT5Base (0.238) -0.02 +0.04 +0.04 +0.05 +0.09
SPLADE+FlanT5XXL (0.472) -0.05 -0.14 +0.12 0.00 -0.02
SPLADE+FlanUL2 (0.545) -0.10 -0.05 -0.04 -0.01 -0.07
Contriever+FlanT5Small (0.318) -0.15 +0.05 -0.05 -0.04 +0.06
Contriever+FlanT5Base (0.302) -0.07 +0.02 +0.01 +0.02 +0.05
Contriever+FlanT5XXL (0.356) 0.00 0.00 +0.12 +0.12 +0.15
Contriever+FlanUL2 (0.455) -0.02 +0.04 +0.09 +0.08 +0.01
Table 5. LaMP-2
Disparity Intervals
BM25+FlanT5Small (0.886) -0.03 +0.01 -0.01 -0.03 -0.01
BM25+FlanT5Base (0.907) -0.02 +0.01 0.00 -0.01 -0.04
BM25+FlanT5XXL (0.859) -0.02 +0.07 +0.01 +0.01 -0.02
BM25+FlanUL2 (0.728) -0.08 -0.05 -0.04 0.00 -0.10
SPLADE+FlanT5Small (0.847) +0.01 +0.05 +0.03 +0.01 +0.04
SPLADE+FlanT5Base (0.902) -0.01 +0.03 0.00 -0.01 -0.02
SPLADE+FlanT5XXL (0.864) -0.02 +0.08 0.00 +0.01 0.00
SPLADE+FlanUL2 (0.682) -0.02 +0.14 -0.07 +0.02 +0.02
Contriever+FlanT5Small (0.876) -0.02 0.00 -0.02 0.00 -0.01
Contriever+FlanT5Base (0.894) 0.00 +0.03 0.00 +0.02 0.00
Contriever+FlanT5XXL (0.865) -0.02 +0.07 +0.01 -0.02 +0.01
Contriever+FlanUL2 (0.680) -0.01 0.00 +0.01 +0.10 +0.04
Table 6. LaMP-3
Disparity Intervals
BM25+FlanT5Small (0.217) -0.06 0.00 +0.02 +0.01 0.00
BM25+FlanT5Base (0.223) -0.06 0.00 +0.03 +0.01 +0.02
BM25+FlanT5XXL (0.322) -0.05 +0.11 +0.03 +0.03 +0.05
BM25+FlanUL2 (0.381) -0.07 +0.06 +0.07 +0.01 +0.04
SPLADE+FlanT5Small (0.235) -0.07 -0.01 +0.02 +0.03 +0.02
SPLADE+FlanT5Base (0.268) -0.10 -0.03 +0.02 0.00 +0.02
SPLADE+FlanT5XXL (0.342) -0.06 +0.09 +0.05 +0.03 +0.04
SPLADE+FlanUL2 (0.429) -0.12 +0.04 +0.02 +0.01 +0.01
Contriever+FlanT5Small (0.254) -0.09 -0.02 0.00 +0.01 0.00
Contriever+FlanT5Base (0.268) -0.10 -0.02 +0.01 0.00 +0.01
Contriever+FlanT5XXL (0.367) -0.09 +0.06 +0.01 +0.01 +0.03
Contriever+FlanUL2 (0.449) -0.15 +0.01 -0.01 0.00 +0.02
Table 7. LaMP-4
Disparity Intervals
BM25+FlanT5Small (0.343) +0.01 +0.12 +0.06 +0.04 +0.06
BM25+FlanT5Base (0.507) -0.04 -0.04 -0.01 -0.02 -0.01
BM25+FlanT5XXL (0.508) -0.03 +0.16 +0.02 +0.02 0.00
BM25+FlanUL2 (0.635) -0.07 +0.06 0.00 +0.01 -0.03
SPLADE+FlanT5Small (0.378) -0.03 +0.05 +0.03 +0.01 +0.02
SPLADE+FlanT5Base (0.470) -0.01 -0.01 +0.01 0.00 +0.03
SPLADE+FlanT5XXL (0.495) -0.02 +0.14 +0.09 +0.03 +0.01
SPLADE+FlanUL2 (0.622) -0.06 +0.04 -0.01 0.00 -0.01
Contriever+FlanT5Small (0.377) -0.03 +0.07 0.00 0.00 +0.03
Contriever+FlanT5Base (0.478) -0.02 +0.03 +0.06 -0.02 +0.03
Contriever+FlanT5XXL (0.496) -0.02 +0.18 +0.04 +0.02 +0.04
Contriever+FlanUL2 (0.641) -0.07 +0.05 +0.01 -0.03 -0.02
Table 8. LaMP-5
Disparity Intervals
BM25+FlanT5Small (0.425) -0.12 -0.07 -0.04 -0.07 -0.03
BM25+FlanT5Base (0.421) -0.09 -0.03 -0.03 -0.04 -0.03
BM25+FlanT5XXL (0.536) -0.03 +0.03 +0.04 -0.01 +0.05
BM25+FlanUL2 (0.580) -0.04 +0.03 +0.06 +0.01 +0.02
SPLADE+FlanT5Small (0.362) -0.06 -0.02 +0.03 +0.02 +0.01
SPLADE+FlanT5Base (0.361) -0.02 +0.02 +0.03 +0.04 +0.05
SPLADE+FlanT5XXL (0.527) -0.02 +0.03 +0.03 +0.03 +0.05
SPLADE+FlanUL2 (0.572) -0.02 +0.03 +0.05 +0.02 +0.07
Contriever+FlanT5Small (0.351) -0.05 -0.02 +0.03 +0.01 +0.04
Contriever+FlanT5Base (0.373) -0.04 +0.01 0.00 +0.04 +0.03
Contriever+FlanT5XXL (0.526) -0.02 +0.01 +0.02 +0.02 +0.06
Contriever+FlanUL2 (0.551) -0.01 +0.04 +0.05 +0.05 +0.08
Table 9. LaMP-6
Disparity Intervals
BM25+FlanT5Small (0.490) -0.02 -0.01 +0.01 +0.03 0.00
BM25+FlanT5Base (0.673) -0.04 -0.04 -0.11 -0.03 -0.05
BM25+FlanT5XXL (0.626) +0.02 -0.05 -0.05 -0.04 +0.03
BM25+FlanUL2 (0.685) +0.04 -0.02 +0.01 +0.02 +0.04
SPLADE+FlanT5Small (0.525) -0.05 -0.07 -0.04 -0.01 -0.03
SPLADE+FlanT5Base (0.659) -0.03 -0.06 -0.06 -0.10 -0.02
SPLADE+FlanT5XXL (0.518) +0.07 +0.06 +0.09 +0.04 +0.05
SPLADE+FlanUL2 (0.665) +0.09 -0.01 +0.04 +0.02 +0.03
Contriever+FlanT5Small (0.440) +0.02 +0.03 +0.05 +0.06 +0.08
Contriever+FlanT5Base (0.580) +0.06 +0.04 0.00 +0.02 +0.06
Contriever+FlanT5XXL (0.607) -0.01 -0.02 -0.03 -0.04 -0.04
Contriever+FlanUL2 (0.721) +0.02 -0.06 -0.02 -0.03 -0.01
Table 10. LaMP-7
Appendix E Data Statistics
E.1. LaMP data statistics for Flan-T5-Small
Dataset #queries Avg # Docs (Std) Avg # Pos Labels (Std) Avg % Pos Labels

### v4

@@ Abstract.
Despite the central role of retrieval in retrieval-augmented generation (RAG) systems, much of the existing research on RAG overlooks the well-established field of fair ranking and fails to account for the interests of all stakeholders involved. In this paper, we conduct the first systematic evaluation of RAG systems that integrate fairness-aware rankings, addressing both ranking fairness and attribution fairness, which ensures equitable exposure of the sources cited in the generated content. Our evaluation focuses on measuring item-side fairness, specifically the fair exposure of relevant items retrieved by RAG systems, and investigates how this fairness impacts both the effectiveness of the systems and the attribution of sources in the generated output that users ultimately see. By experimenting with twelve RAG models across seven distinct tasks, we show that incorporating fairness-aware retrieval often maintains or even enhances both ranking quality and generation quality, countering the common belief that fairness compromises system performance. Additionally, we demonstrate that fair retrieval practices lead to more balanced attribution in the final responses, ensuring that the generator fairly cites the sources it relies on. Our findings underscore the importance of item-side fairness in retrieval and generation, laying the foundation for responsible and equitable RAG systems and guiding future research in fair ranking and attribution.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 2025 International ACM SIGIR Conference on Innovative Concepts and Theories in Information Retrieval (ICTIR); July 18, 2025; Padua, Italybooktitle: Proceedings of the 2025 International ACM SIGIR Conference on Innovative Concepts and Theories in Information Retrieval (ICTIR) (ICTIR ’25), July 18, 2025, Padua, Italydoi: 10.1145/3731120.3744599isbn: 979-8-4007-1861-8/2025/07ccs: Computing methodologies Natural language generationccs: Information systems Evaluation of retrieval results
@@ 3.1. Construction of a Test Collection with Utility Labels
We define an item’s worthiness by the marginal gain in utility (utility-gain) it provides to a language model (specifically, the generator in a RAG system) when used to solve a specific task as part of the augmentation process. To assess this utility-gain, each item in the corpus is individually supplied to the generator along with an input question. The utility-gain is then calculated as the difference between the utility of the augmented generator and that of a baseline language model without any information about the item. Formally, let $u_{i}$ denote the baseline string utility score from the vanilla language model prompted only with the input question, and let $u_{j}$ represent the utility score from the language model with a prompt augmented by the $j$’th item $d_{j}$ in the corpus. The item $d_{j}$ is considered useful if the utility-gain $\delta_{j}=u_{j}-u_{i}$ is positive, and not useful otherwise.
@@ 3.3.1. Expected Exposure in the Context of Machine Users
Expected Exposure (EE) (Diaz et al., 2020) works by estimating the exposure of items across rankings (e.g., $\sigma$) created by a subject model, and comparing them with an optimal set of rankings that always satisfy the item-fairness. To represent the attention over $n$ items in the corpus given by the consumer (generator in RAG), an $n\times 1$ system exposure vector $\mathbf{\epsilon}$ is created. This is then compared with an $n\times 1$ target exposure vector $\mathbf{\epsilon}^{*}$, where it represents the exposure of items allocated by an oracle retriever that always rank useful items above non-useful ones (Diaz et al., 2020).
@@
where $m$ is the number of useful items in the corpus of size $n$.
@@ 3.3.3. Expected Attribution Rate
where $\text{NLI}(d,\hat{y}_{\pi})\!=\!1$ if item $d$ entails the output $\hat{y}_{\pi}$, and 00 otherwise, by a natural language inference model. 333 This metric is similar to the context utilization metric from Ru et al. (2024), which also measures how effectively the generator leverages retrieved context. However, Ru et al. (2024) decompose the generated response into multiple claims and focus only on the proportion of relevant items, whereas our approach applies a single NLI check for each retrieved item.
@@ 4. Experiments
We choose the LaMP benchmark (Salemi et al., 2024b) for our dataset. It assesses the personalization capability of language models through retrieval-augmentation of users’ interaction history in a platform. LaMP includes various prediction tasks, such as classification, regression, and generation, and is well-suited for tasks where multiple items can be relevant/useful, unlike QA tasks with typically one or two provenance items. The retrieval items in LaMP have clear providers and consumers, aligning with our goal to ensure fairness for individual item providers. For example, in LaMP-1, retrieval items are academic papers, where exposure can increase citation counts for authors. In LaMP-4, retrieval items are news articles, where exposure can lead to monetary compensation for journalists. [^4] Throughout this paper, we use LaMP-1 to represent classification and LaMP-4 to represent text generation in visualizations. All LaMP tasks were used in our experiments, and similar trends were observed. Due to the absence of a test set, we constructed a test collection as described in §3.1, using the first thousand entries of a user-based development set. Then, we discarded entries that have only one useful item in the corpus, as it is unnecessary to concern item-fairness in that case. We release the test collection along with dataset statistics.
@@
Figure 3. Effect of a fairness control parameter ($\alpha$) on the disparity of rankings ($\overline{\text{EE-D}}$) in LaMP task 1 and 4. The RAG model is configured with the Contriever and Flan-UL2. Each data point represents the normalized EE-D of each run of the experiment (i.e., one query $\rightarrow$ $N$ sampled rankings $\rightarrow$ $\overline{\text{EE-D}}$ of the $N$ rankings).
@@ 5. Results
Table 2. Each value in the table is the difference between the utility of a baseline (deterministic) RAG model and the average utility of a fairer RAG model at a specific retrieval disparity interval. Nonnegative differences are highlighted. All LaMP tasks shows similar trend. A superscript * denotes a statistically significant difference between the treatment performance and the baseline performance based on an unpaired Student’s t-test ($p<0.05$).
@@
Figure 5. Major relationships between retrieval fairness (RetDisparity), ranking quality, expected attribution rate, and consumption fairness (AttDisparity). We empirically show that there is a tradeoff between retrieval fairness and expected attribution rate as there is a positive relationship between ranking quality and attribution rate. Most importantly, we show that retrieval fairness does not necessarily directly propagate to the consumption fairness due to varying attribution rate of a generator. The same interpolation methods are used as Figure 4 and runs were averaged across generators. RetDisparity refers to the disparity among retrieved items, while AttDisparity refers to the disparity among attributed sources.
@@ 7. Conclusion
In future work, we hope to extend this framework to consider graded or missing judgments and exploring the different notions of fairness in RAG systems, ultimately advancing the field of trustworthy RAG systems research.  [^7] We release our code and dataset at https://github.com/kimdanny/Fair-RAG.
Acknowledgements.
This work was supported by NSF grant 2402874. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect those of the sponsors.
@@ Appendix A Notation
Appendix B Utility Labels Labeling Procedure
@@ C.1. Normalization of EE-D
For upper bound, recall that the $\mathbf{\epsilon}$ is computed based on samples of rankings from a stochastic policy $\pi$. For relevance-independent browsing models (e.g., UM), all rankings $\pi\in S_{n}$ have identical exposure norms ${\left\lVert\mathbf{\epsilon}^{\pi}\right\rVert}_{2}^{2}$, where $\mathbf{\epsilon}^{\pi}$ is the exposure of items for ranking sampled from $\pi$. Then,
@@
With UM, the upper bound of EE-D, ${\left\lVert\bar{\mathbf{\epsilon}}\right\rVert}_{2}^{2}$ becomes $\sum_{i=1}^{n}{\text{UM}(i)}^{2}=k$. Therefore, per query $q=\phi_{q}(x)$, we calculate a normalized EE-D
(16) $\widetilde{\text{EE-D}}_{q}={\left\lVert\mathbf{\epsilon}\right\rVert}_{2}^{2}/k\quad\in[0,1]$
@@ C.2. Normalization of EE-R
With UM, the upper bound of EE-R can be calculated by the combination of equation LABEL:eq:te-rel and LABEL:eq:te-non-rel.
@@
(28) $\widetilde{\text{EE-R}}_{q}=\begin{cases}\langle\mathbf{\epsilon},\mathbf{\epsilon}^{*}\rangle/(m+\frac{(k-m)^{2}}{n-m})&\text{($m\leq k$)}\\ \langle\mathbf{\epsilon},\mathbf{\epsilon}^{*}\rangle/(\frac{k^{2}}{m})&\text{($m>k$)}\end{cases}\in[0,1]$
@@ C.3. Normalization of EU
Recall that the string utility $u_{i}=\mu_{u}(y,\mathcal{G}(\phi_{p}(x,\pi_{i})))$ is a utility with the sampled ranking $\pi_{i}\in\Sigma$. We get a new set of utilities $u^{*}_{i}=\mu_{u}(y,\mathcal{G}(\phi_{p}(x,\pi^{*}_{i})))$, where $\pi^{*}_{i}$ is an element of $\Sigma^{*}$, which is a set of sampled permutations (rankings) from a stochastic oracle retriever $\mathcal{S}^{*}$. This stochastic oracle retriever is a policy that always places relevant items above non-relevant items, thus having $m!(n-m)!$ number of unique optimal permutations.
To get the empirical upper bound of the utility $u_{max}$, we take the maximum of $u$ and $u^{*}$:
(29) $u_{max}=max(u_{1},u_{2},\cdots,u_{N},u^{*}_{1},u^{*}_{2},\cdots,u^{*}_{N})$
With $u_{max}$, we max-normalize EU. Since $\frac{1}{N}\sum^{N}_{i=1}\frac{u_{i}}{u_{max}}$ is the same as $(\frac{1}{N}\sum^{N}_{i=1}u_{i})/u_{max}$, per data instance $x$, we get a normalized EU
(30) $\widetilde{\text{EU}}_{x}=\frac{\text{EU}_{x}}{u_{max}}\quad\in[0,1]$
@@
Appendix D Test Collection Statistics per Model
D.1. LaMP corpus statistics for Flan-T5-Small
Corpus #queries Avg # Docs (Std) Avg # Useful Docs (Std) Avg % Useful Docs